# On the wall-normal velocity variance in canonical wall-bounded turbulence

### v1

@@ 4 Discussion
Besides the observed scaling, Eq. (11) also provides insight into the three main questions of the study given in §1 and highlighted in Fig. 1. These questions are discussed in the subsequent sections: the asymptotic limit of $B_{3}$ and its Reynolds-number dependence are further evaluated in §4.1, the deviations observed for the ZPG-TBL flow are explained in §4.2, and the universality of $B_{3}$ for canonical wall-bounded flows is considered in §4.3.
@@ 4.1 Reynolds-number dependence of wall-normal variances
The finite Reynolds-number dependence in Eq. (13) should theoretically be the same for the streamwise and spanwise variances assuming the spectrum is isotropic at the Kolmogorov scales (Spalart, 1988; Perry & Li, 1990), with the only difference being a lower constant $C_{v}$ for the spanwise $v$ component aligned with the spanwise wavenumber $k_{y}$ (Saddoughi & Veeravalli, 1994).
The revised variance prediction is evaluated for the DNS profiles in Fig. 7(a). The horizontal lines represent $B_{3}$ and the profiles are compensated by $f_{3}(Re_{\epsilon})$ in Eq. (13) to offset the variance “missing” due to the finite Reynolds number. The compensation leads to an observed collapse of the DNS along $B_{3}\approx$ 1.55 in the range $z^{+}=$ 10 to 100, with discrepancies seen only for the cases with $Re_{\tau}=$ 180.
Figure 7: Evaluation of Eq. (11) for the wall-normal variance. (a) The DNS profiles compensated by the local stress $u_{\tau z}$ and the Reynolds number dependence $f_{3}(Re_{\epsilon})$ in Eq. (13), with horizontal lines indicating $B_{3}=$ 1.55 (dotted) $\pm$0.1 (dashed). (b) Variances measured at $z=0.1\delta$ for varying $Re_{\tau}$, where the lines are $B_{3}=$ 1.55 (dotted) $\pm$0.1 (dashed) with the correction $f_{3}(z^{+})$ in Eq. (14). The markers in (b) correspond to literature values: ($\bullet$) present DNS (Sillero et al., 2013; Lee & Moser, 2015; Yao et al., 2023); ($\circ$) pipe DNS (Pirozzoli et al., 2021); ($\Diamond$) channel DNS (Hoyas et al., 2022); ($\triangleright$) pipe experiments (Zhao & Smits, 2007); (++) TBL experiments (Fernholz & Finley, 1996); ($\times$) TBL experiments (De Graaff & Eaton, 2000); (*) TBL experiment (Deshpande et al., 2020).
@@
The same leading-order term $C{z^{+}}^{-1/2}$ has been used in previous spectrum-based estimates for the Reynolds-number dependence (Perry et al., 1986). However, the coefficient $C$ varies significantly across studies, including 0.65 (Perry et al., 1986), 1.55 (present), 4.37 (Spalart, 1988), and 5.58 (Perry & Li, 1990).
The $f_{3}(z^{+})$ relation in Eq. (14) is combined with $B_{3}\approx 1.55\pm 0.1$ in Fig. 7(b) to evaluate the present findings against previous measurements of wall-normal variances. The literature measurements defined in the Fig. 7 caption are specific to $z/\delta=$ 0.1 at the limit of the logarithmic region and are shown relative to $u_{\tau z}$. There is close agreement with $B_{3}-f_{3}(z^{+})$ across the included range of measurements. Most of the points are within the narrow uncertainty range corresponding to variability in large-scale energy seen in Fig. 6(a). The measurements outside of the dashed lines are all from laboratory experiments, which have a higher level of uncertainty due to challenges in measuring $w$ across all scales. A notable exception to the agreement in Fig. 7(b) is CICLoPE pipe measurements (Örlü et al., 2017), which are excluded here because the variances exceed 2 with local scaling and are considered to be an outlier.
@@ 4.2 Wall-normal variances in enclosed and ZPG flows
The previous results and Eq. (11) support the variance $\overline{w^{\prime 2}}\sim u_{\tau z}^{2}$ being proportional to the local stress rather than the surface scale $U_{\tau}^{2}$. This distinction is important when comparing statistics between different flow configurations that can have different stress profiles $\tau(z)/\tau_{w}=u_{\tau z}^{2}/U_{\tau}^{2}$. Example stress profiles are shown in Fig. 8(a). The profiles are at matched Reynolds number $Re_{\tau}\approx$ 2 000 to minimize any differences due to Reynolds-number effects.
Figure 8: Profiles for flow cases at matched Reynolds number $Re_{\tau}\approx$ 2 000. (a) Stress $\tau=\rho u_{\tau z}^{2}$, where channel and pipe cases have the same linear decay. (b) Variances relative to $U_{\tau}^{2}$. (c) Variances relative to $u_{\tau z}^{2}/(1-z/\delta)$.
The channel and pipe flows in Fig. 8(a) both have linearly decaying stresses with $u_{\tau z}^{2}=U_{\tau}^{2}(1-z/\delta)$. In contrast, the ZPG-TBL flow has an approximately sigmoidal shape with $u_{\tau z}\approx U_{\tau}$ near the surface. These differing shapes result directly from the mean momentum balance at $z=0$, which reduces to $dP/dx=d\tau/dz$ with the stress decay balancing the mean pressure gradient. Equation (2) and a constant near-surface stress is therefore specific to ZPG flows. In pressure-driven flows like a pipe or channel there is a non-negligible stress decay throughout the logarithmic region.
The different stress profiles can be compensated using $u_{\tau z}^{2}/(1-z/\delta)$ to normalize the variance. This normalization is identical to $U_{\tau}^{2}$ for a linear stress profile, and for the ZPG flow it compensates for the difference in $u_{\tau z}^{2}$ from the linear decay. The compensated variances in Fig. 8(c) show much closer agreement between the flow configurations than the surface normalization in Figs. 1(a) and 8(b). Compensating for the stress profile also accounts for the peak farther from the surface observed in Fig. 1(b). The higher $w$ variance and peak value farther from the wall can therefore be attributed predominately to the local scaling $\overline{w^{\prime 2}}\sim u_{\tau z}^{2}$ and the impact of the pressure gradient on the near-wall stresses.
The remaining differences between flows in Fig. 8(c) are small but non-negligible. They cannot be attributed to effects of $f_{3}(Re_{\epsilon})$ in Eq. (13) because the $Re_{\epsilon}$ profiles in Fig. 3(c) overlap closely at the matched $Re_{\tau}$. The remaining differences are more likely due to the moderately larger footprint of turbulent eddies at low wavenumbers for the ZPG-TBL flow, which is evident from Figs. 4(b), 5(f), and 6(a). This footprint is discussed in further detail below.
@@ 4.3 Active and inactive motions
The analysis thus far utilized a primitive decomposition of large- and small-scale contributions to the variance shown in Fig. 2(a). For the purpose of this discussion, a more physically meaningful classification of the turbulent eddies is the concept of active and inactive motions (Townsend, 1961). Per Townsend’s orginal hypothesis, at a given height $z_{1}$ in a population of superimposed attached eddies, active motions are centered near $z_{1}$ and contribute to the local Reynolds shear stress $\overline{u^{\prime}w^{\prime}}(z_{1})$. Inactive motions are larger and centered farther from the wall, and the portion of the eddy extending down to $z_{1}$ does not contribute to the correlation between $u$ and $w$ at $z_{1}$ (Townsend, 1961; Deshpande et al., 2021).
@@
Inactive motions contribute substantially to the streamwise and spanwise variance (Deshpande et al., 2021), but the AEH considers $w$ fluctuations to be purely active (Townsend, 1961) with negligible inactive contributions (Bradshaw, 1967; Katul et al., 1996). The wall-normal $w^{\prime}$ fluctuations are thus closely related to the local shear stress, especially for the most energetic scales in the spectrum $E_{ww}$. It is therefore reasonable for these active motions to be proportional to $u_{\tau z}$ corresponding to the local stress. The dominance of active motions can then explain the observed scaling for both the peak in the spectrum and the overall variance. This local scaling is equivalent to the wall scaling in Eq. (2) for ZPG flows in the high-$Re_{\tau}$ limit, but not for enclosed flows with a linear stress decay as seen in Fig. 8(a).
Figure 9: Comparison of Reynolds shear stress spectra $E_{uw}(k_{y})$ at fixed $z$ positions in viscous and outer units.
@@
While $w$ fluctuations are predominately active, recent investigations indicate the presence of inactive contributions to both $w$ and $uw$, highlighting a limitation in Townsend’s hypothesis (Deshpande et al., 2025). The AEH predicts statistics for canonical wall-bounded flow based solely on a linear superposition of fluctuations emerging from a hierarchy of vorticity-dominated structures whose center scales with distance from the wall, i.e. geometrically self-similar eddies. In real wall-bounded flows, however, these attached eddies coexist alongside non-self-similar motions including viscosity-dominated small scales and superstructures or very-large-scale motions, and there is non-linear interaction across all the motions (Guala et al., 2011; McKeon, 2017). These non-linear interactions result in small contributions from the inactive motions to both the $w$ and $uw$ signals (Deshpande et al., 2025).
@@
These inactive contributions can be detected via an energy decomposition methodology (Deshpande et al., 2021) requiring specific multi-point statistics that are not readily available for all the DNS datasets considered herein. The presence of active and inactive $w$ motions are instead inferred here by qualitatively comparing the $E_{ww}$ spectra in Fig. 4 with the $E_{uw}$ shear stress spectra shown in Fig. 9. The latter exhibits the same scaling as $E_{ww}$: the wavenumber and amplitude of the $E_{uw}$ peak are well described by $\ell_{\epsilon}$ and $u_{\tau z}$ and the normalized peak value is approximately 1 at each of the $z$ positions shown. This reaffirms the validity of Townsend’s (1976) hypothesis that the $w$ and $uw$ signals are dominated by contributions from the active (i.e. localized) motions.
@@
While the scaling of the spectral peaks and surrounding wavenumbers is consistent between $E_{ww}$ and $E_{uw}$, the differences in Figs. 4 and 9 before the peak show a clear failure of the local scaling at low wavenumbers. Further, the contribution from the largest spanwise eddies to $E_{ww}(k_{y})$ grows with distance from the wall as seen in Fig. 4, which does not necessarily contribute to the correlation of $u$ and $w$ at that same scale, as can be noted from the trends in Fig. 9. The spectra thus indicate that the contributions to $w$ and $uw$ at the low spanwise wavenumbers are likely coming from the inactive motions. The contribution of these inactive motions to the overall variance, however, is small based on Fig. 6(a). On considering the energy distribution across frequency or streamwise wavenumber $k_{x}$ (not shown here), the spectra $E_{ww}(k_{x})$ and $E_{uw}(k_{x})$ are closely related even at the lowest wavenumber, suggesting the minor contribution of inactive motions is more difficult to discern from the dominant active motions in signals along time or $x$ (Deshpande et al., 2021).
@@
The weak signature of inactive motions in Fig. 4 corresponds with the wavenumbers of large- and very-large-scale motions, owing to which it is stronger in the outer region than in the lower log region. These motions are different for enclosed and non-enclosed flows (Monty et al., 2009), due to the influence of the opposite wall on global modes and the presence of a turbulent-nonturbulent interface in the latter case. The footprint of these motions varies with $z$, the type of flow and its Reynolds number, consistent with the trends seen in Fig. 5(f). It is therefore expected that the inactive motions, which deviate from $u_{\tau z}$ scaling, are responsible for small discrepancies in $B_{3}$ including those seen in Fig. 8(c), thus precluding universality in the constant even at asymptotically high Reynolds numbers. Importantly, the inactive contributions become stronger at higher Reynolds numbers (Deshpande et al., 2025).

### v2

@@ 4 Discussion
Besides the observed dependence on $u_{\tau z}$ and $\ell_{\epsilon}$, Eq. (11) also provides insight into the three main questions of the study given in §1 and highlighted in Fig. 1. These questions are discussed in the subsequent sections: the asymptotic limit of $B_{3}$ and its Reynolds-number dependence are further evaluated in §4.1, the deviations observed for the ZPG-TBL flow are explained in §4.2, and the universality of $B_{3}$ for canonical wall-bounded flows is considered in §4.3.
@@ 4.1 Reynolds-number dependence of wall-normal variances
The finite Reynolds-number dependence in Eq. (13) should theoretically be the same for the streamwise and spanwise variances assuming the spectrum is isotropic at the Kolmogorov scales (Spalart, 1988; Perry and Li, 1990), with the only difference being a lower constant $C_{v}$ for the spanwise $v$ component aligned with the spanwise wavenumber $k_{y}$ (Saddoughi and Veeravalli, 1994).
The revised variance prediction is evaluated for the DNS profiles in Fig. 7(a). The horizontal lines represent $B_{3}$ and the profiles are compensated by $f_{3}(Re_{\epsilon})$ in Eq. (13) to offset the variance “missing” due to the finite Reynolds number. The compensation leads to an observed collapse of the DNS along $B_{3}\approx$ 1.55 in the range $z^{+}=$ 10 to 100, with discrepancies seen only for the cases with $Re_{\tau}=$ 180. Increases above the predicted value begin near $z/\delta\approx$ 0.5 for every case, consistent with the decrease in $\ell_{\epsilon}$ observed in Fig. 3(b) and the trends seen in Figs. 5 and 6.
Figure 7: Evaluation of Eq. (11) for the wall-normal variance. (a) The DNS profiles compensated by the local stress $u_{\tau z}$ and the Reynolds number dependence $f_{3}(Re_{\epsilon})$ in Eq. (13), with horizontal lines indicating $B_{3}=$ 1.55 (dotted) $\pm$0.1 (dashed). (b) Variances measured at $z=0.1\delta$ for varying $Re_{\tau}$, where the lines are $B_{3}=$ 1.55 (dotted) $\pm$0.1 (dashed) with the correction $f_{3}(z^{+})$ in Eq. (14), and the solid line excludes the higher order term. The markers in (b) correspond to literature values: ($\bullet$) present DNS (Sillero et al., 2013; Lee and Moser, 2015; Yao et al., 2023); ($\circ$) pipe DNS (Pirozzoli et al., 2021); ($\Diamond$) channel DNS (Hoyas et al., 2022); ($\triangleleft$) Couette DNS (Lee and Moser, 2018; Hoyas and Oberlack, 2024); ($\triangleright$) pipe experiments (Zhao and Smits, 2007); (++) TBL experiments (Fernholz and Finley, 1996); ($\times$) TBL experiments (De Graaff and Eaton, 2000); (*) TBL experiment (Deshpande et al., 2020).
@@
The same leading-order term $C{z^{+}}^{-1/2}$ has been used in previous spectrum-based estimates for the Reynolds-number dependence (Perry et al., 1986). However, the coefficient $C$ varies significantly across studies, including 0.65 (Perry et al., 1986), 1.55 (present), 4.37 (Spalart, 1988), and 5.58 (Perry and Li, 1990).
The $f_{3}(z^{+})$ relation in Eq. (14) is combined with $B_{3}\approx 1.55\pm 0.1$ in Fig. 7(b) to evaluate the present findings against previous measurements of wall-normal variances. The literature measurements defined in the Fig. 7 caption are specific to $z/\delta=$ 0.1 at the limit of the logarithmic region and are shown relative to $u_{\tau z}$. There is close agreement with $B_{3}-f_{3}(z^{+})$ across the range of measurements, including for turbulent plane Couette flow which is not represented in the Table 1 cases. Most of the points are within the narrow uncertainty range corresponding to variability in large-scale energy seen in Fig. 6(a). The measurements outside of the dashed lines are all from laboratory experiments, which have a higher level of uncertainty due to challenges in measuring $w$ across all scales. A notable exception to the agreement in Fig. 7(b) is CICLoPE pipe measurements (Örlü et al., 2017), which are excluded here because the variances exceed 2 with local scaling and are considered to be an outlier.
To demonstrate the relative importance of the two terms in Eq. (14), the solid line in Fig. 7(b) represents $f_{3}(z^{+})$ without the empirical 3/2-3/2 term. The higher-order correction representing the dissipative cutoff is significant for $Re_{\tau}\lesssim$ 1 000 (i.e. $z^{+}\lesssim$ 100 in the figure), but for $Re_{\tau}\approx$ 2 800 its impact on the total wall-normal variance is less than 1%, beyond which the 1/2-1/2 term representing the inertial subrange is dominant.
@@ 4.2 Wall-normal variances in enclosed and ZPG flows
The previous results and Eq. (11) support the variance $\overline{w^{\prime 2}}\sim u_{\tau z}^{2}$ being proportional to the local stress rather than the surface scale $U_{\tau}^{2}$. This distinction is important when comparing statistics between different flow configurations that can have different stress profiles $\tau(z)/\tau_{w}=u_{\tau z}^{2}/U_{\tau}^{2}$, with example profiles shown in Fig. 8(a). The profiles are at matched Reynolds number $Re_{\tau}\approx$ 2 000 to minimize any differences due to Reynolds-number effects. Fig. 8 includes a turbulent plane Couette flow at the same Reynolds number (Hoyas and Oberlack, 2024). Couette flows are both enclosed and zero-pressure-gradient, thus allowing a direct assessment of factors that may lead to the different profile shapes for different canonical flows in Fig. 1.
Figure 8: Profiles for flow cases at matched Reynolds number $Re_{\tau}\approx$ 2 000 including DNS of a plane Couette flow (Hoyas and Oberlack, 2024). (a) Stress $\tau=\rho u_{\tau z}^{2}$, where channel and pipe cases have the same linear decay. (b) Variances relative to $U_{\tau}^{2}$. (c) Variances relative to $u_{\tau z}^{2}/(1-z/\delta)$.
The channel and pipe flows in Fig. 8(a) both have linearly decaying stresses with $u_{\tau z}^{2}=U_{\tau}^{2}(1-z/\delta)$. In contrast, the ZPG-TBL flow has an approximately sigmoidal shape with $u_{\tau z}\approx U_{\tau}$ near the surface and the Couette flow has a nearly constant stress profile. These differing shapes result directly from the mean momentum balance at $z=0$, which reduces to $dP/dx=d\tau/dz$ with the stress decay balancing the mean pressure gradient. Equation (2) and a constant near-surface stress is therefore specific to ZPG flows including Couette. In pressure-driven flows like a pipe or channel there is a non-negligible stress decay throughout the logarithmic region.
The different stress profiles can be compensated using $u_{\tau z}^{2}/(1-z/\delta)$ to normalize the variance (Tuerke and Jiménez, 2013). This normalization is identical to $U_{\tau}^{2}$ for a linear stress profile, and for the ZPG flows it compensates for the difference in $u_{\tau z}^{2}$ from the linear decay. The compensated variances in Fig. 8(c) show much closer agreement between all flow configurations than the surface normalization in Figs. 1(a) and 8(b). Compensating for the stress profile also accounts for the peak farther from the surface observed in Fig. 1(b). The higher $w$ variance and peak value farther from the wall can therefore be attributed predominately to the local stress dependence $\overline{w^{\prime 2}}\sim u_{\tau z}^{2}$ and the impact of the pressure gradient on the near-wall stresses.
[…]
The remaining differences between flows in Fig. 8(c) are small but non-negligible, most notably the higher ZPG-TBL peak. They cannot be attributed to effects of $f_{3}(Re_{\epsilon})$ in Eq. (13) because the $Re_{\epsilon}$ profiles in Fig. 3(c) overlap closely at the matched $Re_{\tau}$. The remaining differences are more likely due to the moderately larger footprint of turbulent eddies at low wavenumbers for the ZPG-TBL flow, which is evident from Figs. 4(b), 5(f), and 6(a). This footprint is discussed in further detail below.
@@ 4.3 Active and inactive motions
The analysis thus far utilized a primitive decomposition of large- and small-scale contributions to the variance shown in Fig. 2(a). For the purpose of this discussion, a more physically meaningful classification of the turbulent eddies is the concept of active and inactive motions (Townsend, 1961). Per Townsend’s orginal hypothesis, at a given height $z_{1}$ in a population of superimposed attached eddies, active motions are centered near $z_{1}$ and contribute to the local Reynolds shear stress $\overline{u^{\prime}w^{\prime}}(z_{1})$. These motions, consequently, are expected to be directly proportional to the local stress. Inactive motions are larger and centered farther from the wall, and the portion of the eddy extending down to $z_{1}$ does not contribute to the correlation between $u$ and $w$ at $z_{1}$ (Townsend, 1961; Deshpande et al., 2021).
@@
Inactive motions contribute substantially to the streamwise and spanwise variance (Deshpande et al., 2021), but the AEH considers $w$ fluctuations to be purely active (Townsend, 1961) with negligible inactive contributions (Bradshaw, 1967; Katul et al., 1996). The wall-normal $w^{\prime}$ fluctuations are thus closely related to the local shear stress, especially for the most energetic scales in the spectrum $E_{ww}$. It is therefore reasonable for these active motions to be proportional to $u_{\tau z}$ corresponding to the local stress. The dominance of active motions can then explain the observed dependencies for both the peak in the spectrum and the overall variance. This local stress is equivalent to the wall scaling in Eq. (2) for ZPG flows in the high-$Re_{\tau}$ limit, but not for enclosed flows with a linear stress decay as seen in Fig. 8(a).
Figure 9: Spectra as a function of wavenumber $k_{y}$ and wall-normal position $z$ at matched Reynolds number $Re_{\tau}\approx$ 2000 for the channel and ZPG TBL cases. Rows correspond to $E_{ww}$ (top) and the shear stress spectrum $E_{uw}$ (bottom).
@@
In its original form, Townsend’s AEH predicts statistics for canonical wall-bounded flow based solely on a linear superposition of fluctuations, which emerge from a hierarchy of kinematic structures whose center scales with distance from the wall, i.e. geometrically self-similar eddies. In real wall-bounded flows, however, these attached eddies coexist alongside non-self-similar motions including viscosity-dominated small scales and superstructures or very-large-scale motions (VLSMs), and there is non-linear interaction across all the motions (Guala et al., 2011; McKeon, 2017). These non-linear interactions, primarily driven by ‘inactive’ superstructures/VLSMs (Deshpande et al., 2021), lead to small $\delta$-scaled contributions to both the $w$ and $uw$ signals (Deshpande et al., 2025). This highlights a limitation of Townsend’s original hypothesis, whereby inactive motions make a small yet non-negligible contribution to the wall-normal Reynolds-stress components that are considered purely active under Townsend’s AEH. Further discussion on the interaction with inactive motions is given elsewhere (Deshpande et al., 2025).
@@
These inactive contributions can be detected via an energy decomposition methodology (Deshpande et al., 2021) requiring specific multi-point statistics that are not readily available for all the DNS datasets considered herein. The presence of active and inactive $w$ motions are instead inferred here by qualitatively comparing $E_{ww}$ spectra with the $E_{uw}$ shear stress spectra shown in Fig. 9 at matched Reynolds number for the channel and ZPG TBL cases. The peaks in the $E_{ww}$ and $E_{uw}$ spectra occur at similar wavenumbers and exhibit the same wall-normal trends in the lower half of the boundary layer. This similarity reaffirms the validity of Townsend’s (1976) hypothesis that the $w$ and $uw$ signals are dominated by contributions from the active (i.e. localized) motions.
@@
While the scaling of the spectral peaks and surrounding wavenumbers is consistent between $E_{ww}$ and $E_{uw}$, the differences at the lowest wavenumbers for fixed $z/\delta$ positions in Fig. 4(b) show a clear failure of the local stress normalization at the largest scales. The contribution from the largest spanwise eddies to $E_{ww}(k_{y})$ grows with distance from the wall as seen for wavenumbers $k_{y}\delta\lesssim$ 10 in Fig. 9(a,b), but there is an absence of energy at these same wavenumbers in Fig. 9(c,d) such that the largest eddies along $y$ do not necessarily contribute to the correlation of $u$ and $w$ at that same scale. Combined with the failure of the local stress scaling, this suggests that the low spanwise wavenumber contributions to $w$ and $uw$ are likely coming from the inactive motions. The contribution of these inactive motions to the overall variance, however, is small based on Fig. 6(a). On considering the energy distribution across frequency or streamwise wavenumber $k_{x}$ (not shown here), the spectra $E_{ww}(k_{x})$ and $E_{uw}(k_{x})$ are closely related even at the lowest wavenumber, suggesting the minor contribution of inactive motions is more difficult to discern from the dominant active motions in signals along time or $x$ (Deshpande et al., 2021).
@@
The weak signature of inactive motions in Fig. 4 corresponds with the wavenumbers of large- and very-large-scale motions, owing to which it is stronger in the outer region than in the lower log region. These motions are different for enclosed and non-enclosed flows (Monty et al., 2009), due to the influence of the opposite wall on global modes and the presence of a turbulent-nonturbulent interface in the latter case. The footprint of these motions varies with $z$, the type of flow and its Reynolds number, consistent with the trends seen in Fig. 5(f). It is therefore expected that the inactive motions, which deviate from the dependence on $u_{\tau z}$, are responsible for small discrepancies in $B_{3}$ including those seen in Fig. 8(c), thus precluding universality in the constant even at asymptotically high Reynolds numbers. Importantly, the inactive contributions become stronger at higher Reynolds numbers (Deshpande et al., 2025).